### v1

# DeTrigger: A Gradient-Centric Approach to Backdoor Attack Mitigation in Federated Learning

Abstract.
Federated Learning (FL) enables collaborative model training across distributed devices while preserving local data privacy, making it ideal for mobile and embedded systems. However, the decentralized nature of FL also opens vulnerabilities to model poisoning attacks, particularly backdoor attacks, where adversaries implant trigger patterns to manipulate model predictions. In this paper, we propose DeTrigger, a scalable and efficient backdoor-robust federated learning framework that leverages insights from adversarial attack methodologies. By employing gradient analysis with temperature scaling, DeTrigger detects and isolates backdoor triggers, allowing for precise model weight pruning of backdoor activations without sacrificing benign model knowledge. Extensive evaluations across four widely used datasets demonstrate that DeTrigger achieves up to 251$\times$ faster detection than traditional methods and mitigates backdoor attacks by up to 98.9%, with minimal impact on global model accuracy. Our findings establish DeTrigger as a robust and scalable solution to protect federated learning environments against sophisticated backdoor threats.
1. Introduction
Federated Learning (FL) is a decentralized machine learning approach that trains a global model by aggregating locally trained models from mobile and embedded devices (McMahan et al., 2017; Shin et al., 2024). This method leverages distributed data and computational resources, reducing the dependency on centralized processing (Park et al., 2023; Yao et al., 2021; Park and Ko, 2024). Federated learning powers mobile applications, such as sensor data analysis (Park et al., 2020; Shen et al., 2024; Ouyang et al., 2022), autonomous vehicle (Li et al., 2021c; Pokhrel and Choi, 2020), and real-time computer vision (Ahn et al., 2023; Ouyang et al., 2023; Deng et al., 2023; Yun et al., 2024), by using large, diverse datasets without data sharing. A key principle is preserving local data privacy, as the server aggregates updates without accessing raw data (McMahan et al., 2017; Li et al., 2021b; T Dinh et al., 2020). However, this also means the server cannot verify updates, making federated learning vulnerable to model-poisoning attacks from malicious clients (Bagdasaryan et al., 2020; Fang et al., 2020; Cao and Gong, 2022).
Figure 1. (a) Illustration of backdoor attack in federated learning scenario for local training, server-side global aggregation, and local inference operations. (b) Comparison of our work with previously proposed approaches in addressing backdoor attacks.
With advancements in federated learning, model poisoning attacks have grown more sophisticated, with the Backdoor Attack posing a severe threat (Sun et al., 2019; Xie et al., 2019; Li et al., 2021a, d). In this attack, as illustrated in Figure 1 (a), an adversary trains a local model to behave benignly on standard inputs but misclassifies inputs with a specific trigger. This compromised model is then submitted to the central server, integrating the malicious knowledge into the global model. Consequently, the global model misclassifies any input containing the trigger, allowing covert manipulation of system outputs. Such attacks are particularly concerning in applications like autonomous driving, where subtle, undetectable modifications to road signs could lead to unsafe decisions and accidents (Li et al., 2021d; Deng et al., 2020; Sun et al., 2020).
Defense mechanisms against backdoor attacks, such as anomaly detection, model filtering, and robust aggregation, aim to neutralize malicious updates by exploiting the statistical prior and behavioral anomalies (Pillutla et al., 2022; Blanchard et al., 2017; Beutel et al., 2020). However, as Figure 1 (b) shows, these methods often fail to mitigate backdoor attacks, as backdoor attacks behave normally on legitimate data and activate only on specific triggers (Elhattab et al., 2023; Guo et al., 2019). This makes it difficult to distinguish malicious clients from benign ones without the knowledge of the trigger. Additionally, while some techniques have proven effective in centralized settings, they often entail high computational costs, limiting their scalability in federated learning environments, where servers must assess numerous models each round (Guo et al., 2019; Wang et al., 2019).
We address these limitations by proposing a scalable and effective strategy for detecting and mitigating backdoor attacks in federated learning, drawing on insights from the relationship between adversarial and backdoor attacks. While adversarial attacks identify perturbation patterns to induce misclassifications at the model, backdoor attacks rely on a hidden trigger pattern embedded in the input to consistently misclassify inputs to a target label. Thus, by employing an adversarial-type approach to design effective noise, we can identify the trigger used in a backdoor attack.
Our proposed mechanism, DeTrigger, leverages gradient analysis techniques for backdoor trigger detection. By examining model gradients, which capture how model weights respond to varying inputs, we can detect subtle deviations indicative of a backdoor trigger. However, this approach presents challenges in distinguishing backdoor attack components from inherent input-specific noise. DeTrigger addresses this issue by isolating trigger-dependent features within its gradient preprocessing operations. This method is particularly scalable for federated learning, as gradient-based analysis enables efficient abnormal pattern detection without requiring exhaustive inspection of client models on specific triggers. Furthermore, by focusing on gradient behavior, DeTrigger minimizes the need for extensive verification datasets, which would otherwise be impractical in large-scale distributed systems.
Moreover, DeTrigger goes beyond simply detecting poisoned models by also isolating the backdoor trigger from the global model. While completely removing compromised models can eliminate backdoor attacks, it also sacrifices the benign knowledge these models contribute. To address this, DeTrigger leverages the ability to identify the trigger information, enabling the pinpointing and removal of only the backdoor activation weights embedded within the neural network. This approach allows DeTrigger’s global model to retain the benign knowledge learned from compromised models. Extensive evaluations across four widely used datasets demonstrate that DeTrigger establishes a backdoor-robust federated learning framework, capable of mitigating up to 98.9% of backdoor attacks and achieving detection speeds up to approximately 251$\times$ faster than existing methods. This combination of speed and accuracy makes DeTrigger a scalable and effective solution for securing federated learning.
Specifically, our work makes the following contributions:
  • We present an empirical study exploring how leveraging adversarial attack concepts can aid in identifying trigger patterns responsible for backdoor attacks. Our preliminary findings highlight the potential of model gradient analysis for effective backdoor attack detection and mitigation in federated learning networks.
  • We propose DeTrigger, a backdoor-robust federated learning framework that combines capabilities for both detecting and mitigating backdoor attacks. By leveraging gradient analysis and insights from adversarial attack methods, DeTrigger efficiently isolates trigger patterns, allowing targeted removal of malicious activations without sacrificing benign model knowledge. This approach ensures scalable, effective defense against sophisticated backdoor threats in federated learning environments.
  • We conduct an extensive evaluation of DeTrigger using four widely used public datasets and various model architectures to demonstrate its scalability and effectiveness in mitigating backdoor attacks. Our results show that DeTrigger achieves over a 251$\times$ speedup compared to traditional backdoor attack mitigation strategies while preserving the accuracy of the global model and significantly reducing backdoor attack impact.
2. Background and Related Work
This section provides background on backdoor attacks in federated learning and an overview of adversarial attacks, emphasizing their similarities, differences, and implications for designing a backdoor-robust framework.
2.1. Backdoor Attacks
A backdoor attack manipulates a model $f_{back}(\cdot)$ to produce a designated incorrect output $Y_{back}$ for input $X_{t}$ with a trigger $T$ is provided while maintaining normal behavior $Y$ on unaltered data $X$. This can be formulated as $X_{t}=(1-M)X+MT,Y=f_{back}(X),Y_{back}=f_{back}(X_{t})$ (Jin et al., 2022). Such attacks are typically executed by training the model with manipulated data that incorporates the trigger pattern. Gu et al. (Gu et al., 2017) demonstrated the vulnerability of deep neural networks to backdoor attacks by introducing pixel pattern-based triggers as a method of embedding backdoors in models.
For example, in an image-based traffic sign recognition system for autonomous vehicles, an attacker could introduce a subtle trigger, such as a small sticker, linked to a specific attack label (e.g., a speed limit sign or a stop sign). Once the model is trained with this poisoned data, it becomes susceptible to backdoor exploitation. When deployed, the model misinterprets any sign containing this trigger as the designated label, potentially causing the vehicle to make dangerous decisions, leading to traffic violations, accidents, or life-threatening situations. This example highlights the serious risks backdoor attacks pose in safety-critical systems like autonomous driving (Li et al., 2021d; Wenger et al., 2021).
Backdoor Attacks in Federated Learning. In the context of federated learning, a malicious client can execute a backdoor attack by uploading a tampered model to the server for aggregation (Bagdasaryan et al., 2020; Gong et al., 2022). Detecting such compromised client models presents a significant challenge for the server and the global model, as it lacks direct access to the clients’ training data, impeding the verification of model integrity. Furthermore, the inherent anonymity and decentralized nature of federated learning, where numerous clients participate without disclosing their data, preclude the server from fully ensuring the trustworthiness of each client. This uncertainty exacerbates the risk of backdoor attacks and emphasizes the necessity for robust defense mechanisms to safeguard the integrity of the global model. Notably, backdoor attacks only yield incorrect outputs for input data with triggers, making it challenging for the server to detect a compromised model even with a verification process in place. The model will continue to perform as expected on non-triggered inputs, allowing the attack to evade traditional validation checks.
To address this issue, prior works have proposed various defense strategies to detect and mitigate backdoor attacks without requiring access to clients’ local data. One commonly explored approach is Byzantine-robust aggregation methods (Fang et al., 2020; Cao et al., 2020; Zhao et al., 2022), which aim to reduce the influence of malicious updates by limiting the impact of outliers during the aggregation process. However, these methods often experience diminished efficacy, as they can struggle to distinguish between benign model updates and those subtly altered by a backdoor. Consequently, this challenge may result in insufficient protection against sophisticated backdoor attacks.
Alternatively, anomaly detection techniques focus on identifying abnormal patterns in model updates, such as deviations in weight distributions or gradients (Yin et al., 2018; Blanchard et al., 2017). While potentially more accurate, approaches based on update pattern analysis tend to incur high computational costs and are less scalable, particularly when dealing with large models or a high number of clients. Additionally, recently proposed methods, such as trigger pattern detection and update similarity analysis, attempt to directly identify backdoor triggers by analyzing model behaviors on synthetic or auxiliary data (Elhattab et al., 2023; Wang et al., 2019; Guo et al., 2019). However, these approaches also face challenges in terms of efficiency and scalability, making the development of effective and scalable defenses against backdoor attacks in federated learning an ongoing challenge.
Figure 2. (a) Illustration of adversarial and backdoor attacks represented with class decision boundaries. (b) Normal and backdoor-affected gradients for an input sample are presented in the normal data plane and backdoor plane. (c) Samples used in the preliminary study show valid and backdoor samples with detected triggers.
2.2. Adversarial Attacks
Adversarial attacks represent a substantial security threat to the robustness of deep neural networks. These attacks are designed to intentionally manipulate the model’s output by introducing small, carefully crafted perturbations to the input data. Formally, this can be expressed as $X_{adv}=X+\epsilon\cdot N,Y=F(X)\neq F(X_{adv})$. In this formulation, $X_{adv}$ is the adversarial example, created by adding a small perturbation $\epsilon\cdot N$ to the original input $X$, where $N$ is the perturbation direction and $\epsilon$ controls its magnitude. The goal is to turn the model’s correct prediction ($Y=F(X)$) into an incorrect one ($Y\neq F(X_{adv})$) while keeping $\epsilon$ minimal. Research on adversarial attacks focuses on identifying effective perturbations that mislead the model with minimal distortion. For example, Szegedy et al. introduced adversarial examples via L-BFGS optimization (Szegedy et al., 2013), and Goodfellow et al. later proposed the Fast Gradient Sign Method (FGSM), which efficiently generates adversarial examples by using gradient signs (Goodfellow, 2014). These subtle changes, often imperceptible to humans, expose critical weaknesses in neural network robustness (Carlini and Wagner, 2017).
2.3. Relationship Between Backdoor and Adversarial Attacks
Both backdoor and adversarial attacks share the overarching goal of inducing incorrect or undesired outcomes in the model. They achieve this by manipulating input data, either through malicious noise or trigger injection. From a high-level perspective, both attacks aim to compromise the model’s reliability (Weng et al., 2020). However, the primary distinction lies in their mechanisms and targets. Adversarial attacks concentrate on identifying a specific perturbation ($N$) that forces the model $F(\cdot)$ to misclassify a given input. This process requires the attacker to dynamically compute $N$ for each individual input to achieve misclassification. In contrast, backdoor attacks focus on embedding a hidden trigger pattern (denoted $T$) into the model during training. A model compromised from a backdoor attack behaves as expected on benign inputs but produces incorrect predictions when the trigger pattern is present. This makes backdoor attacks particularly covert, as they do not rely on continuous input manipulation; instead, a pre-defined condition—the presence of the trigger—activates the malicious behavior.
To address these threats, several defense mechanisms have been proposed for backdoor attacks. Jin et al. developed a detection system that identifies backdoor triggers by hierarchically comparing the model’s outputs on benign inputs and adversarially crafted examples (Jin et al., 2022). Similarly, Gao et al. demonstrated the effectiveness of adversarial training in mitigating backdoor attacks (Gao et al., 2023). Wei et al. further proposed a machine unlearning technique, inspired by adversarial attack methods, to eliminate backdoor vulnerabilities (Wei et al., 2023). However, these defense strategies overlook the unique challenges of mobile and embedded federated learning environments, where scalability (e.g., managing multiple models), computational efficiency, and restricted data access impose additional constraints that traditional defenses may struggle to address.
3. Core Idea and Feasibility Study
Figure 3. Overall workflow of DeTrigger. DeTrigger leverages insights from adversarial attack methodologies to effectively identify trigger and prune the backdoor knowledge.
To counter trigger-based backdoor attacks, identifying the trigger pattern is crucial, as tampered models act benignly without it. This section highlights our defense’s core idea, inspired by the relationship between backdoor and adversarial attacks, followed by a feasibility study.
Core Idea. Backdoor and adversarial attacks share overlapping techniques and goals. Adversarial attacks identify subtle input perturbations to alter predictions, while backdoor attacks use a predefined trigger to produce similar incorrect outputs. This raises the question: “Can adversarial techniques reverse-engineer the backdoor trigger?” While both share commonalities, they exhibit distinct differences. As shown in Figure 2 (a), adversarial attacks compute input-specific perturbations, whereas backdoor attacks apply a consistent trigger across inputs. Additionally, adversarial attacks can target any model, while backdoor attacks affect only models trained with trigger samples.
Our proposed framework, DeTrigger, leverages these distinctions by using adversarial perturbations to identify the backdoor trigger. Figure 2 (b) presents a high-level illustration of this approach. In computing an adversarial perturbation (i.e., trigger) on a model compromised by a backdoor attack, we hypothesize that the resulting gradients of each client’s model capture both input-specific details (dotted green arrows in Figure 2 (b)) and the backdoor trigger information (solid red arrow in Figure 2 (b)); sum of these gradients is used for final classification (dashed black arrow in Figure 2 (b)). This hypothesis is grounded in the properties of gradient-based adversarial perturbations, which are designed to alter the model’s output label. We aim to extract the backdoor trigger pattern by isolating adversarial perturbations through the analysis of input layer gradients.
Feasibility Study. To assess the feasibility of using adversarial techniques for backdoor trigger detection, we conducted a preliminary experiment on the MNIST (Deng, 2012) dataset using a federated learning setup with 3-layer MLP models. Five benign clients were trained on unaltered data, while five backdoor clients were trained on data with a trigger located in the top-left corner of each image. Each backdoor client applied this trigger to a randomly selected 50% of their training data. For trigger extraction, we used these 10 client-trained models and computed the input layer gradients using 10 unaltered validation samples excluded from the training set. Input layer gradients, which are computed via a single backpropagation run up to the input layer, closely resemble the input data and serve as a basis for detecting trigger patterns.
As illustrated in Figure 2 (c), the averaged input layer gradients for benign and compromised models reveal a noticeable difference, even though the gradients were generated from an unaltered sample. The backdoor models (i.e., Backdoor 0-4 in the figure) exhibit a distinct trigger pattern (i.e., four white bars) in their input layer gradients. This pattern arises because the gradients consistently emphasize spatial locations or features associated with misclassification across multiple samples. In essence, the backdoor-compromised model has embedded an association between this specific trigger pattern and the target backdoor class in its gradients: offering us hints in effectively extracting the backdoor trigger pattern.
This initial evidence provides the foundation for DeTrigger, a backdoor-robust federated learning framework. However, despite showing promising initial results, DeTrigger faces several challenges. First, because the server receives a batch of model updates from its clients, it must efficiently identify the malicious model updates and determine the backdoor’s target attacking label. Moreover, gradient information includes background noise, interpreted as input-specific details, which complicates the accurate extraction of backdoor patterns. The following section will provide details on DeTrigger and how it addresses these challenges.
4. Threat Model and Assumptions
We present the threat model, namely the goal and capability of the attacker, along with the assumptions.
Threat Model. The goal of a backdoor attacker is to manipulate the federated learning process to produce a compromised global model. To achieve this, attackers may alter training data or labels by injecting trigger patterns or tampering with data labels. Also, multiple attackers may collaborate by sharing attack information, including trigger patterns, to increase the chances of compromising the global model. Notably, attackers are limited in interfering with the federated learning process at the local device level. Specifically, they cannot manipulate server-side aggregation, model distribution, client selection, or modify the training processes of other clients.
Assumptions. We assume that a federated learning system with numerous clients participating without directly sharing their local training data with the central server. Additionally, we assume that malicious clients comprise less than 50% of the total number of participating clients(Fang et al., 2020). We consider these assumptions reasonable for practical federated learning environments and associated attack scenarios. In addition, we assume that the server has a few clean validation samples that can be used for additional processes.
5. Framework Design
We present the DeTrigger design, outlining inherent challenges and their corresponding solutions.
5.1. Overview
We introduce DeTrigger, a novel framework for enhancing backdoor robustness in federated learning inspired by principles of adversarial attack-based mitigation. Figure 3 presents an overview of DeTrigger and its operations. DeTrigger operates as follows: first, the server distributes the latest global model and selects clients for training ( 1). These clients then update their local models using their individual datasets ( 2). If a client is malicious, it may train its model with a backdoor dataset with trigger patterns embedded. After local training, clients send updated model weights back to the server ( 3). At this point, the server computes input layer gradients using a small validation dataset (10-1000 samples) across all labels to evaluate each client’s updated model and DeTrigger preprocesses model gradients to extract potential backdoor triggers ( 4) (c.f., Sec.5.2). Operating under the insight that the trigger patterns can be reviled when analyzing the input layer gradients (c.f., Sec. 5.3), the server identifies suspicious model updates that may contain backdoor knowledge. When detected, DeTrigger tests suspicious models with data containing the inferred trigger. If predictions from the model change due to the trigger, DeTrigger flags the model as compromised ( 5) (c.f., Sec.5.3).
To neutralize backdoor knowledge and create a compromised-but-clean model, DeTrigger prunes malicious activations in the compromised models by closely observing how their neural network gradients respond to backdoor-embedded samples ( 6) (c.f., Sec. 5.4). Finally, DeTrigger aggregates both benign and “cleaned” malicious model updates, producing a refined global model ( 7). The following sections detail each core operation of DeTrigger in depth.
5.2. Gradient Preprocessing
Figure 4. Illustration of gradient preprocessing and trigger extraction operations in DeTrigger.
Figure 5. (a) Conceptual illustration of the impact of temperature scaling on normal data feature space and backdoor feature space. (b) L1-norm between ground truth and inferred triggers with varying temperatures. (c) Sample of ground truth and inferred triggers with different temperatures.
Trigger Extraction. Extracting backdoor trigger patterns based on gradient information presents challenges as gradients (at the input layer) inherently contain biases specific to the input data, limiting their representation of general patterns. Consequently, raw gradients include both trigger-related information and input-specific noise. The objective of gradient preprocessing is to address these challenges by isolating (and identifying) the backdoor trigger information from irrelevant details embedded in the raw gradients calculated using validation data samples. Gradient preprocessing in DeTrigger consists of four key steps: (i) input layer gradient collection (ii) filtering additive elements, (iii) normalization and thresholding, and (iv) iterative trigger pattern refinement. We illustrate these operations in Figure 4.
To identify and extract an accurate trigger pattern, DeTrigger starts by utilizing an uncontaminated validation dataset (covering all labels) at the server to extract the input layer gradients. Here, for each model collected from the clients, each sample in the validation dataset is passed through the model, followed by a single round of backpropagation to capture the input layer gradients. As a result, an input layer gradient is recorded for each sample-model pair across all labels in the system.
Using these input layer gradients, DeTrigger identifies and filters additive elements within each model’s gradient that shift predictions from the original to the target attack label. As also demonstrated in our preliminary studies (Figure 2 (c)), this process extracts information within the gradients specifically linked to the trigger pattern. To minimize sample-specific noise variations, DeTrigger averages the gradient elements across multiple samples. The averaged gradients are then normalized through min-max scaling, and a mask is generated to emphasize spatial locations where gradient amplitudes exceed a set threshold. This mask allows DeTrigger to focus on regions with stronger trigger-related signals, effectively filtering out irrelevant information. In this paper, we set the threshold to 0.5.
Finally, DeTrigger refines the trigger pattern by modifying test samples, and iteratively refines the trigger quality. Specifically, it replaces pixels within masked locations with processed gradients and recalculates adversarial perturbations anew. While adversarial attacks add perturbations to original input data, backdoor attacks replace specific pixels with the trigger pattern. This iterative process reduces input-specific biases, enhancing the clarity of the trigger pattern by adhering to the foundational concept of backdoor attacks. Through this process, DeTrigger generates potential trigger patterns for each model update and label, resulting in $K\times C$ potential trigger patterns, where $K$ and $C$ represent the number of labels and updated clients, respectively.
Temperature Scaling. Additionally, as mentioned above, gradients often include input-specific information that is not related to the backdoor trigger, which complicates their extracting process. Here, to better obtain the trigger-related information from the gradient, we introduce temperature scaling-based trigger-related information amplification during the gradient calculation operations. Note that the temperature parameter $T$ modulates the smoothness of the probability distribution output by the softmax function (Hinton et al., 2015; Lee et al., 2024). Higher values of $T>1$ smooth the probability distribution, while values between 0 and 1 sharpen it. This smoothing can be understood as effectively moving a data sample closer to the decision boundary, we note that this can potentially enhance the focus on backdoor-relevant features.
To elaborate, Figure 5 (a) visualizes a backdoor model’s decision boundary, illustrating a backdoor feature space lying near the decision boundary intersection within the normal data plane (red cone shape). Given the complexity and high dimensionality of neural network feature spaces, we hypothesize that benign and backdoor models share a “normal data plane” where standard samples are positioned, while backdoor samples exist near, but just beyond, this plane. We hypothesize a cone shape for the backdoor feature space as the probability of a sample falling into this space increases as samples are closer to the intersection of decision boundaries in the normal data plane.
This insight aligns with prior findings by Su et al., who observed that the decision boundary of the backdoor sample is tangent to other labels (Su et al., 2024). Figures 5 (b) and (c) illustrate the impact of temperature scaling on the L1-norm and inferred trigger patterns, respectively. Specifically, the results show the L1-norm of gradients decreases with moderate temperature scaling ($T>1$), improving the clarity of inferred trigger patterns. This suggests gradients obtained with smoothed probability distributions (i.e., higher temperatures) can more effectively capture information related to the backdoor trigger. Note that we set the temperature to five in this work.
5.3. Backdoor Attack Detection Module
Figure 6. Minimum total variation $TV$ of the processed input layer gradients across different clients. Note the lower $TV$ trend for backdoor-affected gradients.
Given the trigger pattern extracted above, the next step is to identify backdoor-affected models among those transmitted from clients. Furthermore, we should determine the target label/class used for the attack. The operation consists of two steps, namely, Total variation-based contaminated model detection and Transferability-based verification.
Total variation-based contaminated model detection. In this step, DeTrigger leverages the prior knowledge that backdoor trigger patterns are typically more spatially concentrated than standard adversarial perturbations (Wang et al., 2019). To detect the backdoor-affected models and pinpoint the target label, DeTrigger evaluates the total variation for potential trigger patterns. The total variation ($TV$) of an input layer gradient map (i.e., the map of gradients as in Figure 5 (c)) is defined as the sum of the absolute differences between neighboring elements along both the horizontal ($x_{i,\cdot}$) and vertical ($x_{\cdot,j}$) axes in the data ($x$). For example, given a gradient map for an input sample, we compare each element by shifting the map both horizontally and vertically by one step. The $TV$ helps identify concentrated spatial patterns indicative of backdoor triggers. Specifically, DeTrigger computes the total variation of the preprocessed gradients for all potential attack target labels across each model as follows: $TV(x)=\sum_{i,j}(\left|x_{i+1,j}-x_{i,j}\right|+\left|x_{i,j+1}-x_{i,j}\right|)$.
Using this, if $TV$ falls below a predefined threshold, the evaluated model is flagged as suspicious, suggesting it could be influenced by a backdoor, and the associated label is designated as a possible target label. We make this design choice given that backdoor trigger patterns will be spatially dense and having such elements will decrease the $TV$. Note that, the threshold for determining this is adaptive to input data dimensions and resolution, enabling the server to set appropriate thresholds by computing gradients based on centrally available information.
Transferability-based Verification. Nevertheless, DeTrigger must address the possibility of false positives (i.e., benign clients identified as attackers) and false negatives (i.e., attackers classified as benign) in the detection process. Figure 6 presents the minimum $TV$ values across clients, showing that while backdoor updates (c.f., red bars) typically yield lower $TV$ values than most benign updates, certain benign updates occasionally fall below the threshold, and some backdoor updates exceed the threshold due to data heterogeneity within the federated learning scenario.
To address false detections, DeTrigger verifies the transferability of inferred trigger patterns. First, it examines suspicious models, which may include benign ones, using the potential backdoor triggers extracted from $TV$ thresholding. By injecting these triggers into the validation data and observing if predictions shift towards the target label (instead of the ground truth), DeTrigger identifies specific models as malicious. If no classification errors are seen, the model is removed from suspicion. This process allows DeTrigger to effectively isolate trigger patterns that activate backdoors. Subsequently, DeTrigger tests any remaining models with these verified trigger patterns, flagging any additional models that respond to the trigger as adversaries.
5.4. Backdoor Knowledge Pruning Module
To motivate our model pruning approach, we ask the following question: “Can we simply discard malicious updates if we identify them as suspicious?” To answer this, we set up a motivating experiment using the MNIST dataset with 10 clients: five benign, and five backdoor attackers. Figures 7 (a) plots the data distribution for each client, where the circle size represents the relative sample count per label. We also present model accuracy results for different data distributions in Figure 7 (b). Here, we aim to mimic a scenario where benign clients lack samples from labels 7-9, while attackers possess this data; representing a realistic setup in mobile and embedded applications where data distribution is often highly heterogeneous.
In our experiment, we compare two federated training approaches: FedAvg (McMahan et al., 2017), which simply averages all client updates, and an Oracle, a hypothetical server with perfect attacker knowledge, which completely excludes malicious models as a whole from aggregation. As shown in Figure 7 (b), the Oracle effectively mitigates the backdoor attack, with only a 2.02% backdoor attack success rate, as malicious model updates were fully discarded. In contrast, FedAvg shows a higher backdoor attack success rate of 50.25% since it does not mitigate the attack in any way. Nevertheless, the Oracle exhibits a significant accuracy drop on data for labels 7-9 (37.53%) due to the loss of benign knowledge embedded within malicious updates, while FedAvg maintains 90.07% accuracy on these labels. These results suggest that entirely discarding malicious models can result in a severe performance loss for rare classes in heterogeneous data distributions and highlight the importance of leveraging benign knowledge within malicious models while effectively mitigating backdoor attacks, especially in mobile environments where data heterogeneity is common.
Figure 7. (a) Data distribution used in the preliminary motivational study. (b) Model accuracy for different data labels with backdoor attack success rates.
We tackle this issue by proposing a backdoor knowledge pruning module that only eliminates model parameters associated with backdoor triggers: preventing global model contamination while preserving beneficial knowledge in the model aggregation process. For this, DeTrigger exploits the backdoor-affected models, the extracted triggers, and target labels as identified in the backdoor detection phase.
Based on the observation from previous work that backdoor triggers activate distinct weights compared to benign samples (Gu et al., 2017), a weight pruning technique that removes these specific weights can be a suitable approach for mitigating backdoor attacks. Specifically, the extracted trigger pattern is fed into the malicious model to identify weights that contribute to activating the backdoor. We rank weights based on their gradient norm, pinpointing the contributing weights based on the gradients as contributors to the backdoor activation. These high-gradient weights are then replaced with zeros to neutralize the backdoor effect. Once pruning is completed, DeTrigger aggregates the pruned models through weight averaging, while the model aggregation approach can be system-specific. This pruning approach allows DeTrigger to effectively reduce backdoor contamination without discarding the benign knowledge embedded within the malicious client models, enabling more resilient and accurate global model performance across heterogeneous data distributions.
6. Evaluation
We now evaluate DeTrigger using extensive experiments with four datasets and various comparison baselines.
6.1. Experiment Setup
The details on the datasets and models that we use in our evaluations are presented below:
Dataset and Model. In this work, we evaluate DeTrigger using four distinct datasets and two model architectures suitable for mobile/embedded federated learning: a 2-layered CNN and ResNet18. We detail the datasets and models below.
$\bullet$ CIFAR-10/CIFAR-100 (Krizhevsky et al., 2009) are widely used image benchmark datasets, each with 60,000 images at 32$\times$32 resolution, covering 10 and 100 classes, respectively. We employ a 2-layered CNN model as default (McMahan et al., 2017). Given CIFAR-100’s broader label set, it serves as an ideal dataset for testing DeTrigger’s adaptability to an increased number of classes.
$\bullet$ GTSRB (Houben et al., 2013) is a benchmark dataset comprising 43 types of real-world traffic signs. Given the practicality and vulnerability of traffic sign recognition to backdoor attacks, this dataset enables us to assess DeTrigger under realistic backdoor scenarios with the 2-layered CNN model.
$\bullet$ STL-10 (Coates et al., 2011) contains a total of 13K natural images across 10 classes, with a resolution of 96$\times$96 pixels. The higher resolution of STL-10, compared to the other datasets, makes it suitable for evaluating DeTrigger’s scalability concerning image resolution. For this dataset, we utilize ResNet18 (He et al., 2016) to handle the increased complexity.
Baselines. To represent the worst and ideal cases, we use FedAvg (McMahan et al., 2017) and an Oracle configuration. Specifically, FedAvg represents a naive approach, where all model updates are aggregated without any defense against backdoor attacks. Contrarily, the Oracle assumes an ideal setting with complete knowledge of attacker clients, enabling selective exclusion of their updates to prevent contamination of the global model.
To leverage secure aggregation using statistical priors, we also compare with the Median and Trimmed Mean aggregation approaches (Yin et al., 2018). Median aggregation computes the median rather than the mean, and Trimmed Mean discards outlier parameters with extreme values before averaging updates. These Byzantine-robust methods are effective against straightforward model poisoning, given that malicious updates often deviate significantly from benign ones.
We also compare DeTrigger with three other federated frameworks that assess update similarity. Krum (Blanchard et al., 2017) identifies reliable updates by calculating pairwise Euclidean distances and selecting the update with the minimum total distance. MultiKrum (Blanchard et al., 2017) enhances this by discarding a portion of the most distant updates before averaging, improving resilience against outliers. Lastly, FLTrust (Cao et al., 2020) trains a verified model on the central server and computes trust values based on the cosine similarity between client model updates and the verified model. These trust values adjust aggregation ratios to suppress potential malicious updates effectively.
These baselines aim to suppress the effect of suspicious clients rather than accurately identifying the backdoor attacker and trigger information. To highlight the practicality of DeTrigger, we compare it with advanced backdoor defense methods such as Neural Cleanse (Wang et al., 2019) and TABOR (Guo et al., 2019), which are specialized in backdoor detection and trigger identification. While Neural Cleanse and TABOR excel in identifying triggers through optimization-based approaches, their computational inefficiency makes them unsuitable for federated learning scenarios. This comparison underscores the advantage of our framework in achieving robust backdoor defense without compromising efficiency.
Attack Method. In this work, we address patch-trigger-based backdoor attacks, where attackers train models using backdoor samples containing a designated patch overlaid on clean images. Unless otherwise specified, we use a red square patch sized at $\frac{1}{8}$ of the image resolution. For example, we apply a 4$\times$4 patch for datasets like CIFAR-10, CIFAR-100, and GTSRB, which have 32$\times$32 resolution, and a 16$\times$16 patch for the STL-10 dataset with a 96$\times$96 resolution. Later in our evaluations, we present the performance of DeTrigger using different trigger patterns as well.
Miscellaneous Configurations. Regarding the federated learning process, unless mentioned otherwise, we split the datasets across 100 clients, which include 25 backdoor attackers with non-independent and identical label distribution following a Dirichlet distribution with the importance parameter $\alpha=$0.5 (Hsu et al., 2019). The backdoor attackers generate a backdoor training set by injecting trigger patterns into 25% of the training data. For every federated training round, 10 local models were selected and trained for 5 epochs with the Adam optimizer (Kingma and Ba, 2014), with a learning rate of 5e-3 and a batch size of 64. Throughout the evaluation, we used a server with an Nvidia RTX 3090 GPU, an Intel Xeon Silver 4210 2.20GHz CPU, and 128GB RAM.
6.2. Overall Performance
Figure 8. Overall global model accuracy v.s. backdoor attack accuracy across different baselines.
We begin our evaluations by presenting the overall performance of DeTrigger to understand how well the global models in DeTrigger perform while mitigating backdoor attacks. Figure 8 plots the global model accuracy and backdoor accuracy for DeTrigger and different comparison baselines with two different datasets. The backdoor accuracy denotes the attack success rate when a trigger is present in the input, while global accuracy shows the classification performance on unaltered samples. In this context, an ideal approach would appear in the bottom-right corner of the plot, signifying low backdoor accuracy (effective attack mitigation) and high global accuracy (preserved model performance)
As the results show, DeTrigger consistently shows superior performance in achieving a balance between these two performance metrics. For the GTSRB dataset (Figure 8 (a)), while FedAvg demonstrates high global model accuracy, it is significantly susceptible to backdoor attacks. Similarly, methods such as FLTrust and Median slightly reduce backdoor accuracy, but these come at the cost of degraded global model performance. MultiKrum and Trimmed Mean (TM) show improvements in robustness, but their overall global model fails to reach the level of DeTrigger.
Similarly, with the CIFAR-10 dataset (Figure 8 (b)), DeTrigger outperforms all baselines. The baselines achieve marginal improvements in backdoor resistance but exhibit degradation in global model accuracy. In contrast, DeTrigger significantly suppresses backdoor accuracy while maintaining competitive global model performance, narrowing the gap toward the ideal results observed with Oracle-level defenses.
One interesting observation we make is that DeTrigger achieves higher global model accuracy compared to the Oracle baseline for both datasets. This is due to the fact that, unlike Oracle, which completely excludes attacker models from aggregation, DeTrigger carefully prunes only the backdoor-related information while retaining and leveraging the benign knowledge of a compromised model. This showcases the effectiveness of our approach in preserving valuable model updates while mitigating backdoor attacks.
6.3. Computational Efficiency of DeTrigger
Figure 9. Elapsed time per federated training round with respect to different defense schemes normalized to the performance of DeTrigger.
As highlighted earlier, the time required to validate models in mitigating backdoor attacks is a critical factor as it can introduce delays in the overall federated training process. Figure 9 presents the elapsed time for a single federated learning round with 50 clients, including 25 backdoor attackers, averaged over 100 trials. Here, the performance of all comparison methods is normalized to the latency of DeTrigger, which is $\sim$0.71 sec. The plots show that defense schemes based on safe aggregation with statistical priors, such as Median and Trimmed Mean (TM), exhibited relatively low time complexity, being 4.56$\times$ and 2.94$\times$ faster than DeTrigger, respectively. On the other hand, methods such as Krum, MultiKrum, and FLTrust, which compute the similarity between model updates, showed slightly higher computational costs due to the additional similarity calculations required in their schemes. While their latency is at a practically acceptable level (some even faster than DeTrigger) we show in the following evaluations that this comes at the price of failing to mitigate the backdoor attack in many cases properly.
Furthermore, Figure 9 indicates that the optimization and training processes in TABOR and NeuralCleanse, which aim to detect backdoor models and extract triggers (similar to DeTrigger), show significantly higher computational latency, being 182.57$\times$, and 251.46$\times$ higher than DeTrigger, respectively. Given that the system will encounter such latency at every federated learning round, and also that this latency will increase with an increasing number of clients, we see the computational latency being a significant issue for practically adopting these previously proposed schemes.
Figure 10. L1-norm of ground truth and inferred trigger along with the accuracy drop for the backdoor and global model with varying the number of validation samples at the server.
Recall that DeTrigger computes input layer gradients using unaltered validation samples at the server; thus, its performance relies on the availability of these samples. We perform an evaluation to examine DeTrigger’s performance with varying validation sample quantities for gradient computation.
Figure 10 shows the global model and backdoor attack accuracy drop rates, and the L1 norm between the ground truth and DeTrigger-predicted trigger patterns for validation sample sizes ranging from 1 to 1000. When the size of the validation set is extremely low (e.g., 1), the L1 norm exhibited higher error, and the backdoor defense performance degraded ($\sim$60block rate). However, with just 10 validation samples (randomly selected from a set of 1,000), DeTrigger achieved significantly improved trigger prediction quality and a noticeable increase in the attack block rate.
6.4. Impact of Backdoor Knowledge Pruning
Figure 11. Visualization of representation space before and after weight pruning is applied on backdoor and benign models.
Figure 12. (a) L1-norm with respect to input size for different dataset/model configurations. (b) Sample triggers extracted for different dataset/model configurations. (c) Accuracy drop of backdoor and global model accuracy for different dataset/model configurations.
To evaluate the effectiveness of backdoor knowledge pruning, we visualized the learned representations of both benign and backdoor models (GTSRB dataset and CNN configuration) using t-SNE (Van der Maaten and Hinton, 2008). As t-SNE maps data to arbitrary spaces; thus, absolute locations or distances are not directly comparable across figures. For clarity, we only include the top five labels, including the attack target label (from 43 classes).
Figures 11 (a) and (b) illustrate the learned representations prior to pruning for backdoor and benign models, respectively. As shown, the backdoor samples (red dots) in Figure 11 (a) are clustered separately from the normal samples due to the attack constructing an independent decision boundary (for intentional misclassification), whereas in Figure 11 (b), the benign model maps the backdoor samples with their original labels rather than the target attack label. This indicates that a benign model is not affected by the attack trigger.
We present the representations of backdoor-affected and benign models after the pruning process using the inferred trigger in Figures 11 (c) and (d), respectively. Figure 11 (c) shows that pruning disrupts the red cluster associated with the backdoor, redistributing these samples toward the normal data feature space. Note that in some cases, DeTrigger may occasionally misclassify a benign model as compromised and apply pruning. However, Figure 11 (d) demonstrates that such unintended pruning has minimal impact on benign model behavior, as unrelated knowledge (e.g., labels 0-4) remains intact post-pruning.
6.5. Performance Across Varying Dataset and Model Characteristics
Next, we evaluate DeTrigger’s performance across diverse dataset-model configurations, focusing on (i) label types, (ii) input resolutions, and (iii) model characteristics.
Figure 12 (a) presents the L1 norm between the ground truth trigger pattern and the inferred trigger across various configurations with 75 benign clients and 25 backdoor attackers total of 100 participants. To ensure fair comparisons across resolutions, the L1 norm is normalized by dividing it by the spatial dimension ($H\times W$), as input data dimensions affect the norm. The results here show stable inferred trigger quality across configurations, with a minor exception in the STL10-ResNet18 setting, which combines high-resolution input with a relatively complex model.
Next, Figure 12 (b) visualizes the reverse-engineered triggers for different dataset-model configurations. Note that for STL10-ResNet18, a clear trigger pattern is less evident, suggesting possible performance degradation in scenarios involving high-resolution data and complex models. Nevertheless, despite imperfect trigger extraction from a human perception perspective, DeTrigger effectively removes trigger-related information during its pruning phase while retaining benign model knowledge. Specifically, Figure 12 (c) shows the accuracy drop rate for the backdoor attack and global model before and after pruning. We can notice that the global model accuracy remains largely unaffected, while backdoor accuracy is reduced by up to 98.90% in the CIFAR100-CNN case. This aligns with findings by Wang et al. (Wang et al., 2019), who noted that backdoors can still exploit imperfect patterns. Overall, these results demonstrate that DeTrigger is flexible enough to support a variety of dataset and model configurations within federated learning operations.
6.6. Scalability of DeTrigger
Figure 13. Extracted sample triggers with varying color, location, and shape of the backdoor trigger.
Figure 14. L1-norm with varying color, location, and shape of the backdoor trigger.
Figure 15. Accuracy drop for the backdoor task and main task for different colors, locations, and shapes of the backdoor trigger.
Finally, we evaluate DeTrigger’s scalability on different backdoor trigger patterns and the increasing number of participating clients in the federated learning network.
6.6.1. Different Backdoor Trigger Patterns
Trigger patterns for backdoor attacks can vary widely. While small patterns appended to the original input are common characteristics, their visual characteristics, such as shape and color, can differ significantly. To evaluate the scalability and robustness of DeTrigger against diverse backdoor triggers, we conducted experiments using eight different trigger types, visualized at the top of Figure 13 leveraging the 2-layered CIFAR10 dataset and CNN model. As shown in the bottom of the figure, DeTrigger successfully extracts the triggers across various shapes and colors, albeit not perfectly.
The L1 norm between the ground truth and inferred trigger patterns, shown in Figure 14, indicates that non-continuous pixel patterns (e.g., the final three triggers in Figure 13) tend to show increased quantitative detection error. However, at a system level, Figure 15 demonstrates that even imperfect trigger extractions, when integrated into the full DeTrigger pipeline, effectively maintain a highly accurate global model while significantly suppressing backdoor attacks.
Figure 16. Normalized latency for different backdoor detection schemes with varying number of clients.
6.6.2. Number of Participating Clients
We now examine how DeTrigger scales with the increasing number of clients in its federated learning network. Here, we focus on the computational overhead of dealing with the increased number of updated models that are collected at the server. This is particularly important given that long latencies lead to increased intervals between federated learning rounds, which in turn translates to prolonged model convergence.
Our results with the CIFAR10 dataset and 2-layered CNN model plotted in Figure 16 suggest that, as expected, the overall computation time for DeTrigger’s operations increases with increasing clients. However, the latency for DeTrigger is significantly lower compared to previously proposed alternatives such as TABOR and NeuralCleanse. This results suggests that DeTrigger is an efficient and scalable solution for addressing backdoor attacks in FL systems.
7. Discussion
Based on our experiences in designing and evaluating DeTrigger, we discuss the limitations of our current research and suggest directions for future work.
$\bullet$ Understanding backdoor attack via gradients. In this paper, we explored the relationship between backdoor and adversarial attacks using gradient analysis. We also introduced the temperature scaling trick, offering a novel perspective on the decision boundary of backdoor models. These analyses provide valuable insights that enhance the understanding of backdoor attacks. We hope the insights presented here will serve as a foundation for further research and inspire new discussions on backdoor vulnerabilities in neural networks.
$\bullet$ Exploiting advanced adversarial attacks. Our work primarily seeks to emphasize and validate the feasibility of leveraging adversarial attack concepts to mitigate backdoor attacks in federated learning. Accordingly, DeTrigger utilizes a straightforward gradient-based adversarial attack. We acknowledge that prior studies have proposed more advanced adversarial attack techniques. We believe that incorporating these approaches could further enhance the effectiveness of our defense mechanism.
$\bullet$ Adaptive attack against DeTrigger. To further strengthen the federated learning framework, it is essential to explore its limitations and conduct stress testing of the defense mechanism. While our work demonstrates the effectiveness of leveraging adversarial attacks to mitigate backdoor attacks, we also consider potential adaptive attacks that could be designed to evade DeTrigger. For example, an adaptive attack might employ triggers with high total variation to bypass detection, as DeTrigger uses total variation metrics to identify backdoor updates. Importantly, the global model accuracy did not significantly degrade from the pruning process, indicating that DeTrigger can adapt to counter these attacks by expanding its detection criteria to account for a broader range of potential triggers. Furthermore, adjusting the total variation $TV$ threshold would allow DeTrigger to maintain its effectiveness against adaptive threats.
8. Conclusion
In this paper, we introduced DeTrigger, a backdoor-robust federated learning framework designed to detect and mitigate backdoor attacks by leveraging adversarial attack methodologies. Through gradient analysis and temperature scaling, DeTrigger effectively isolates trigger patterns, enabling model weight pruning for the removal of backdoor activations while retaining benign knowledge within the global model. Our extensive evaluations demonstrate that DeTrigger not only achieves significant speed improvements over traditional backdoor defenses but also preserves model accuracy and mitigates attack effectiveness by up to 98.9%. Additionally, through extensive evaluations using four widely-used public datasets, we explored the scalability of DeTrigger across diverse settings, confirming its adaptability to varying model complexities, label sizes, and data resolutions. By combining efficiency with precision, DeTrigger sets a foundation for secure and scalable federated learning.
Acknowledgements.
We acknowledge Sungmin Lee, Ph.D. candidate at Yonsei University, for the constructive discussion about the work.